OpenStudy (anonymous):

simplify (10/ x-1)/ (1/2 - 7/x+1)

OpenStudy (anonymous):

This is what you wrote: \[\frac{10}{\left(\frac{3}{2}-\frac{7}{x}\right) (-1+x)} \]Is that correct. If not you need some more parenthesis.

OpenStudy (anonymous):

never mind i just got it

OpenStudy (anonymous):

Excellent! Have a good day.

hero (hero):

@robtobey, he probably meant this \[\frac{10}{x - 1} \div \left(\frac{1}{2} - \frac{7}{x + 1}\right)\]

hero (hero):

@Ko-Ki, how do you know you have it? What did you get for it?

OpenStudy (anonymous):

@Hero Thanks for the comment.

OpenStudy (anonymous):

I got 20/x-13 since. If a complex fraction is really just a fraction divided by another fraction, then perform the division. Then division problem by rewriting as multiplication, factoring, and cancelling.